Carlos makes baskets with Bags of Coffee or Boxes of tea. He has 72 bags of coffee and 56 boxes of tea. He wants to put a equal number of bags of coffee or boxes of tea in each basket using all the coffee and tea. Each basket contains only coffee or tea.
What is the largest number of baskets that he can make ?

How many bags of coffee or boxes of tea will he use in each box?

Bags of Coffee?
Boxes of tea ?

Answer by ikleyn(52782)

.

    72 = 8*9

    56 = 8*7


He should place 9 bags of coffee and 7 boxes of tea to each basket.


The number of baskets is 8.


It is the largest number of baskets that he can make under given conditions.

Solved.

---------

My post-solution note.

    In my view, the problem formulation is not so accurate as it should be.

    From the text, I can not understand, if the baskets for coffee and tea are separate or not.

    My solution and my answer are for the case when EACH basket contain both coffee and tea.


    If you meant separate baskets, then the answer is 8 + 8 = 16.

I just used to think,  that persons who compose the problem for this forum,
are not able  IN  PRINCIPLE  to do it accurately.